# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ VIII-E Summary and Prospects
In this section, a channel model simulator called BUPTCMCCCMG-IMT2030 was introduced, including the design framework, user guidelines, and application examples for 6G potential technologies. To summarize, BUPTCMCCCMG-IMT2030 is a channel model simulator developed for 6G wireless communication by the authors. In the current version, BUPTCMCCCMG-IMT2030 supports simulations for the standardized channel model defined in ITU-R M.2412 [42] and adds a series of extensions for 6G features, such as THz, E-MIMO, ISAC, RIS, and SAGIN. Specifically, BUPTCMCCCMG-IMT2030 supports channel simulation scenarios including UMa, UMi, RMa, InH, and SG in SAGIN, and also supports frequency bands including centimeter wave, mmWave, and THz at 132 GHz. In addition, new channel characteristics, such as the SnS of the multiantenna channel and the sparsity of the THz channel, are also included. Simulation frameworks for RIS-assisted and ISAC channels are also implemented. At present, the 2.0 version of BUPTCMCCCMG-IMT2030 has been released on [260]. In future versions, more scenarios (especially more space and air scenarios for SAGIN), frequency bands (more frequency points for THz), and channel characteristics of potential 6G technologies, will be added based on channel measurement and modeling studies to make BUPTCMCCCMG-IMT2030 a powerful assistance tool for channel model simulation.
IX Conclusions
@@
ITU’s current work on the framework and overall objectives represents the official launch of 6G standardization. The technical evaluation work in the standardization is scheduled for 2024, where the 6G channel research is necessary and critical. A survey of 6G channel measurements and modeling and a tutorial of the 6G channel simulator BUPTCMCCCMG-IMT2030 are presented. The representative requirements and challenges of 6G channel research are summarized first. In the survey, we discussed channel measurements and modeling of potential key technologies, including THz, E-MIMO, ISAC, RIS, and SAGIN. Firstly, for channel measurement, a measurement platform with high accuracy and supporting multiple scenarios is urgently needed. Secondly, for channel modeling, new channel characteristics must be properly considered. Geometry-based stochastic modeling is one of the main modeling methods of 6G, but the model needs to be adjusted under different technologies and scenarios. Finally, in the tutorial, we guided the ITU/3GPP GBSM-based 6G channel simulator, BUPTCMCCCMG-IMT2030, which supports the simulation of the aforementioned various 6G potential technologies. The channel simulator has good backward compatibility and is expected to include more scenarios, frequency bands, and channel characteristics in the future.
@@ Acknowledgments
This research is supported in part by National Science Fund for Distinguished Young Scholars (61925102), National Natural Science Foundation of China (92167202, 62101069, 62201086, 62201087), and Beijing University of Posts and Telecommunications-China Mobile Research Institute Joint Innovation Center.

### v4

@@ VIII-E Summary and Prospects
In this section, a summary of 6G channel simulators that support potential enabling technologies is provided. The BUPTCMCCCMG-IMT2030 channel simulator was subsequently introduced, including the design framework, user guidelines, and application examples for potential 6G technologies. To summarize, BUPTCMCCCMG-IMT2030 is a channel model simulator developed for 6G wireless communication by the authors. In the current version, BUPTCMCCCMG-IMT2030 supports simulations for the standardized channel model defined in ITU-R M.2412 [46] and adds a series of extensions for 6G features, such as ISAC, XL-MIMO, THz, RIS, and SAGIN. Specifically, BUPTCMCCCMG-IMT2030 supports channel simulation scenarios, including UMa, UMi, RMa, InH, and SG in SAGIN, and supports frequency bands, including centimeter waves, mmWaves, and THz waves at 132 GHz. In addition, new channel characteristics, such as the SnS of the multiantenna channel and the sparsity of the THz channel, are also included. Simulation frameworks for the RIS-assisted and ISAC channels are also implemented. At present, the 2.0 version of BUPTCMCCCMG-IMT2030 has been released at [314].
The ISAC and RIS channel modeling methods employ a concatenated approach. For instance, consider a single-segment channel comprising 20 clusters, each containing 20 sub-paths. This concatenated structure results in a total of 20×20=400 clusters and 400×400=160000 paths. Each multipath incorporates parameters such as delay, power, angle, and Doppler, which introduces significant challenges for channel impulse response computation and data storage. Consequently, the simulation complexity escalates from $O(N)$ to $O(N^{2})$. Concurrently, in XL-MIMO technology, the number of antenna elements significantly increases. Characterizing both near-field and spatially non-stationary characteristics further elevates the demand for simulation resources. To achieve an optimal trade-off between computational complexity and simulation fidelity, future research should focus on in-depth investigations into cluster pruning techniques for concatenated channels and the development of simplified concatenated methodologies. Additionally, exploring simplified algorithms for capturing the near-field and spatial non-stationary properties of XL-MIMO, as well as AI-assisted complexity reduction methods, represents promising directions for future investigations. In future versions, more scenarios (especially more space and air scenarios for SAGINs), frequency bands (more frequency points for THz systems), and channel characteristics of potential 6G technologies will be added on the basis of channel measurement and modeling studies to make BUPTCMCCCMG-IMT2030 a powerful assistance tool for channel model simulation.
IX Open Issues on 6G Channel
IX-A Channel Model Standardization for 6G
International standardization organizations have launched focused initiatives to advance channel model development in support of emerging 6G requirements. The 3GPP technical specification group (TSG) for radio access network (RAN) approved two new study items at the RAN #102 meeting: “channel modeling for ISAC” [315], and “channel modeling enhancements for 7–24 GHz” [316]. Technical discussions commenced at the RAN1 #116 meeting and concluded at the RAN1 #121 meeting, spanning a total of nine meetings.
For the ISAC channel model study item, typical sensing targets of UAVs, humans indoors, humans outdoors, automotive vehicles, automated guided vehicles, and objects creating hazards on roads/railways are approved as a starting point [317]. The common framework of the ISAC channel model is composed of a path component of the target channel and a component of the background channel. Several key new features of the ISAC channel have been discussed during the 3GPP standardization process [318, 319]: (1) the RCS, which characterizes the target scattering effect and should be modeled for typical sensing targets; (2) the target channel, which is modeled through the concatenated sub-channels of Tx-target and target-Rx; (3) the background channel, where the bistatic mode can reuse existing communication TRs, while the modeling of the monostatic background channel is approved through multi-reference points; and (4) additional features such as spatial consistency, EO, Doppler, and blockage.
For the 7–24 GHz channel model study item, several key enhancements have been introduced [320]. Four additional modeling components have been incorporated [321]: (1) near-field propagation, which accounts for spherical wavefront effects in array-based simulations; (2) SnS, which enables the modeling of power variations across antenna elements at both the BS and UE sides; (3) cluster number variability, which reflects the sparsity of high-frequency channels and captures the observation that the number of clusters can vary with factors such as environment, frequency, bandwidth, and spatial resolution; and (4) polarization power variability. In addition, large-scale and small-scale parameters have been updated based on newly collected FR3 measurement data, and a new suburban macro (SMa) scenario has been defined with a complete set of parameters [322].
Moreover, several standardization challenges remain. Additional technologies such as RIS and SAGIN require continued investigation to support their integration into standardized channel modeling frameworks. For example, further efforts are needed to address micro-Doppler in ISAC channel modeling, which is a key parameter to be used for detecting and tracking the target. Modeling RCS for different materials, distances, and frequencies is also essential for building a standardized RCS database to support identification.
IX-B AI-enabled Methods in Channel Research
AI has rapidly advanced in recent years and plays a vital role in wireless channel research. Early studies applied machine learning (ML) methods such as support vector machines (SVMs), relevance vector machines (RVMs), and principal component analysis (PCA) to tasks such as channel parameter extraction [323], multipath clustering [324], and scenario classification [325], primarily to improve the automation and robustness of channel modeling. With the rise of deep learning (DL) techniques, models can now extract high-dimensional, multiscale features from large-scale datasets and support accurate predictions of channel parameters, including channel state information, path loss, beam indices, and blockage conditions, by leveraging generative adversarial networks (GANs) [326], deep reinforcement learning [327, 328], and other carefully designed DL networks [329]. Furthermore, with the increasing use of environmental sensing technologies, AI models are now being developed to learn direct mappings between physical environments and wireless propagation behavior. In this context, we propose the concept of a digital twin channel (DTC) [330, 331], which leverages sensing technologies to obtain deterministic wireless environmental information and integrates AI models to generate real-time channel fading states. This approach lays the foundation for proactive and reliable communication decision-making and establishes a new online paradigm that actively addresses channel uncertainty.
The DTC framework holds promise for enabling real-time, environment-intelligent channel modeling and adaptive communication system operation. However, further improvements are needed in terms of accuracy, complexity, and generalizability. First, improving modeling accuracy will require the integration of fine-grained environmental representations and the use of high-resolution sensing data to capture subtle variations in environmental dynamics. Second, addressing computational complexity involves exploring lightweight model architectures, knowledge distillation techniques, and efficient inference strategies that support real-time performance. Third, to enhance generalizability, future models should be designed to adapt to a wide range of environments, mobility patterns, and previously unseen deployment scenarios. This may be achieved through the use of large-scale foundation models that are pretrained on diverse environments [332]. These large channel models are capable of capturing common propagation patterns and can transfer knowledge across scenarios.
IX-C Practical Channel Model-enabled System Performance Analysis
The practical channel model is a key factor in system performance analysis. The existing channel models can effectively support the performance analysis of 5G and previous communication systems. However, for 6G, the technology is developing towards directions such as XL-MIMO and ISAC, which brings new characteristics to the channel. Therefore, the channel model needs to be updated. The channel model proposed in this paper has these features and can support various system performance analyses.
Unlike the existing base station antenna arrays, the extremely large-scale antenna array used in XL-MIMO has near-field and spatial non-stationary characteristics, and has been widely observed [104, 112, 105, 106, 107, 108, 114, 115, 117, 118, 119, 120, 103, 111, 31, 121, 122]. Therefore, it is necessary to analyze various performance changes caused by these new characteristics, such as beam training [333, 334, 335, 336, 337, 338], achievable rate [339, 340, 341, 342, 343] and degrees of freedom [344, 345, 346, 347]. However, traditional channel models cannot characterize the new channel characteristics, making it difficult to support the performance analysis of XL-MIMO. [118] based on the measured results, near-field parameters and spatial non-stationary parameters are introduced to represent the channel, and a near-field spatial non-stationary channel model is established. [338] provided examples on how spherical wavefront and spatially non-stationarity affect the XL-MIMO system performance in terms of beamforming gain and achievable rate. [343] focused on the theoretical analysis and empirical study of near-field MIMO capacity. Based on this foundation, more research directions, including how to consider the actual multipath channel, the spatial non-stationarity of the hybrid field channel in narrowband and wideband systems for near-field technology design, and how to design new beamforming in wideband hybrid field channels to overcome the beam splitting effect, can be further explored in future research.
ISAC, as a novel technology for 6G, relies heavily on accurate channel modeling for system performance and technical evaluation. First, 3GPP and numerous academic studies have investigated the RCS characteristics of typical targets such as UAVs, humans, and vehicles. Precise RCS modeling is critical for target classification, recognition, and localization, especially in complex environments [348]. While traditional positioning depends on LoS paths, NLoS signals—leveraging single reflections with AoA and ToA—can also be exploited in cluttered scenarios [349, 350]. Modeling EOs helps enhance positioning via NLoS paths. Micro-Doppler modeling can assist in recognizing the dynamic behavior and posture of sensing targets, such as small movements of human limb motion [351] or UAV rotor rotation [352]. Moreover, accurate modeling of environmental effects improves clutter suppression and the extraction of target features (e.g., delay, and angle), thus enhancing localization robustness under multipath interference [353, 354]. Finally, modeling the sharing feature of communication and sensing supports mutual optimization and performance evaluation, fostering efficient low-power ISAC system design [355, 356].
Similar to XL-MIMO, RIS deployments often operate in near-field conditions due to large physical apertures, making near-field propagation modeling essential for accurate system performance evaluation. Prior analyses under far-field assumptions suggest that the end-to-end RIS gain scales quadratically with the number of elements ($N^{2}$) [357, 358], potentially outperforming MIMO systems with the same array size. However, recent studies incorporating spherical wave models show that near-field effects fundamentally limit such scaling, and passive RIS cannot surpass traditional MIMO under practical conditions [359]. In addition to near-field effects, existing RIS performance analyses often rely on simplified assumptions, such as Rayleigh/Rician fading, rich scattering, continuous phase shifts, and ideal reflection. These oversimplifications can significantly overestimate system performance, as practical environments may exhibit limited angular spread, discrete phase control, and non-ideal electromagnetic responses. Therefore, future RIS-related performance evaluations should consider: (i) near-field propagation, (ii) realistic angular and fading models, (iii) hardware constraints such as discrete phase control, and (iv) electromagnetic characteristics of RIS elements. These aspects are critical for bridging the gap between theoretical modeling and real-world deployment.
X Conclusions
@@
The ITU’s current work on the framework and overall objectives represents the official launch of 6G standardization. With the progression of standardization efforts, research on channel modeling has become increasingly critical. This paper presents a comprehensive survey of 6G channel requirements, field measurements, modeling methodologies and a released simulator. First, the representative requirements and challenges of 6G channel research are summarized. We discuss channel measurements and modeling of potential key technologies, including ISAC, XL-MIMO, mid-band, THz, RIS, and SAGIN. For channel measurement, a measurement platform with high accuracy that can support multiple scenarios is urgently needed. For channel modeling, new channel characteristics must be properly considered. Geometry-based stochastic modeling is one of the main modeling methods of 6G, but the model needs to be adjusted under different technologies and scenarios. Future research prospects are discussed for each technology. A summary of the released 6G channel simulators is presented below. The BUPTCMCCCMG-IMT2030, developed on the basis of the ITU/3GPP GBSM framework, is introduced. It provides practical support for simulating a wide range of 6G technologies, ensures backward compatibility and is being extended to accommodate additional frequency bands, deployment scenarios, and channel characteristics. Finally, we address emerging perspectives in 6G channel research, including recent 3GPP activities on modeling for 7–24 GHz and ISAC, AI-enabled methods in channel research, and system performance analysis.
@@ Acknowledgments
This research is supported in part by National Natural Science Foundation of China under Grant (62525101), National Natural Science Foundation of China (92167202, 62101069, 62201086, 62201087), and Beijing University of Posts and Telecommunications-China Mobile Research Institute Joint Innovation Center.